OpenStudy (anonymous):

Solve the equation for y. 8x – 9y = 11 Question 2.2. Determine which ordered pair in the form (a, b) is a solution of the given equation. 7a – 5b = 28 (Points : 1) (–3, –2) (–2, –3) (0, 4) (4, 0) Question 3.3. Determine whether the ordered pairs in the form (x, y) are solutions of the given equation and select the correct answer. (3, –8), (4, 4) Yes, both are solutions. No, neither is a solution. The first is a solution, but the second is not.

OpenStudy (amoodarya):

8x – 9y = 11 -9y=11-8x divide by -9 y=-11/9 +8/11 x

OpenStudy (anonymous):

wha??? is that for the first one

OpenStudy (amoodarya):

yeah

OpenStudy (anonymous):

ok its 3x- yover 4=11 for # 3

OpenStudy (nurali):

3x - y/4= 11 To determine if the ordered pair satisfies the equation simply plug the x coordinate into the equation wherever you see and plug the y coordinate into the equation wherever you see y. So let's take the first ordered pair: (3,-8) The x value=3 and the y value= -8. So now wherever you see an x, instead put 3 and wherever you see y, instead put -8. When you do you get: [(3) x (3)] - [(-8) / (4)] = 11. When you simplify you get: 9 + 2= 11 and that does work. So (3, -8) is a solution to this equation. Do the same thing with the second ordered pair. And when you do you should get: [(3) x (4)] -[(4) / (4)]= 11 And once you simplify you get 12-1= 11. So both ordered pair satisfy the equation. Yes, both are solutions.

OpenStudy (haseeb96):

Q.1 8x-9y=11 8x-11=9y 8x-11\9=y or y=8x-11\9 Q.2 solution (4,0) 7a-5b=28 a=4 and b=0 now put it on the given equation 7(4)-5(0)=28 28-0=28 28=28 hence answer is correct

OpenStudy (haseeb96):

no, neither is a solution it is correct 100%